OpenStudy (anonymous):

The diameter of a basketball rim is 18 inches. A standard basketball has a circumference of 30 inches. How much space is between the ball and the rim if someone makes a shot and the ball goes exactly through the center of the rim?

OpenStudy (lgbasallote):

how much space means you need to fnd the volume so find the volume of a basketball with the 18-inch diameter and the volume of a standard basketball...then subtract them do you need more help?@jayda

OpenStudy (lgbasallote):

@jayda

OpenStudy (anonymous):

so what would the diameter of the standard basketball be

OpenStudy (lgbasallote):

circumference = \(\pi d\) diameter = \(\frac{circumference}{\pi}\)

OpenStudy (lgbasallote):

so divide 30 by pi to get the diameter

OpenStudy (anonymous):

so 9.6.

OpenStudy (lgbasallote):

yup that's the diameter

OpenStudy (anonymous):

and is the formula 4/3 x pi x r squared?

OpenStudy (lgbasallote):

r cube \[\large V = \frac{4}{3} \pi r^3\]

OpenStudy (anonymous):

woops yeah i meant cubed

OpenStudy (lgbasallote):

so can you solve the two volumes now?

OpenStudy (anonymous):

yeah

OpenStudy (lgbasallote):

great! so to summarize.... it's \[\huge V_{\text{rim}} - V_{\text{basketball}} = \text{answer}\]

OpenStudy (lgbasallote):

V means volume ;)

OpenStudy (anonymous):

652?

OpenStudy (lgbasallote):

hmm let me check..

OpenStudy (anonymous):

i dont think its right. its none of my answers. ha

OpenStudy (lgbasallote):

\[V_{rim} = \frac{4}{3} \pi (9)^3\] \[V_{rim} = 3054\] \[V_{ball} = \frac{4}{3} \pi (4.8)^3\] \[V_{ball} = 463\] were those what you got?

OpenStudy (anonymous):

no i didnt get them.. but i tried it again and i got them. idk why i didnt get them the first time? ha

OpenStudy (lgbasallote):

hmm maybe you didnt divide by 2 to get the radius?

OpenStudy (anonymous):

im not sure. so then subtract 3054 and 463?